OpenStudy (anonymous):

Is this correct? 25^2+10^2=c^2 625+100=c^2 725=c^2 √725=√c^2 √25√29=√c^2 5√29=c

OpenStudy (anonymous):

This is the question. Find the value of C. Express the answer as a simplified radical. http://static.k12.com/eli/bb/811/7537/0/2_36614_44211/7537/a5f1016a929c7017dcd86e01b372da4fa290c833/media/5b84520eeed4d482d0fa592efc2269a0585e47e4/mediaasset_649786_1.gif

OpenStudy (kendricklamar2014):

Let's solve your equation step-by-step. 252+102=c2 Step 1: Simplify both sides of the equation. 725=c2

OpenStudy (kendricklamar2014):

Step 2: Subtract c^2 from both sides. 725−c2=c2−c2 −c2+725=0

OpenStudy (kendricklamar2014):

Step 3: Subtract 725 from both sides. −c2+725−725=0−725 −c2=−725

OpenStudy (kendricklamar2014):

Step 4: Divide both sides by -1. −c2/−1= −725/−1 c2=725

OpenStudy (kendricklamar2014):

Step 5: Take square root. c=±sqrt(725) c=26.925824,−26.925824

OpenStudy (kendricklamar2014):

Answer: c=26.925824,−26.925824

OpenStudy (anonymous):

But it's supposed to be a simplified radical

OpenStudy (isaiah.feynman):

Sure its right.

OpenStudy (anonymous):

I'm right?

OpenStudy (kendricklamar2014):

yes

OpenStudy (anonymous):

Oh ok. Thanks